Solve the triangle, given a=10,b=13, and angle A=25 degrees

3. Solve the triangle, given a = 10, b = 13, and ZA = 25°. (round answer to one decimal place.
Answer #1

Solution: Given theet B 10 A 25 13 MLA In AABC m2A-25° a =10 b = 13 claim: To find the side C and angle B and C By using Law250 + 33.26 it mLC - 180 58.26°+mk (=180° 6 mLC: 180-58.26 O mLC = 121,7 Letratio ① 13 10 sin25o e sin 121.70 sin 33.3° consi